OpenStudy (anonymous):

Physics stuff!! A ball is thrown upward with a speed of 40 feet per second from the edge of a cliff 500 feet above the ground. What is the speed of the ball when it hits the ground? Use acceleration due to gravity as –32 feet per second squared and approximate your answer to 3 decimal places.

OpenStudy (anonymous):

@ganeshie8 @nincompoop @pooja195 @Preetha any ideas?

OpenStudy (empty):

I believe there are two ways to do this 1) Calculate the trajectory 2) Calculate the energy change

ganeshie8 (ganeshie8):

presuming the ball lands on the ground and not on the cliff... even though it was thrown vertically upwards

OpenStudy (anonymous):

this is a calc I problem by the way

ganeshie8 (ganeshie8):

maybe the angle shouldn't matter here because the ball gains its speed by the time it reaches cliff. You may start with the given acceleration due to gravity : \[v(t)-v(0)=\int\limits_0^t a(t)\,dt \]

OpenStudy (anonymous):

okay s v(0) = 40 but how do we find v(t) ?

OpenStudy (empty):

Continue further, \[v(t) = \frac{d x(t)}{dt}\]

OpenStudy (anonymous):

wait slow down , how did you find s(t) ?

ganeshie8 (ganeshie8):

\[\begin{align}v(t)-v(0)&=\int\limits_0^t a(t)\,dt \\~\\ &=\int\limits_0^t -32\,dt \end{align}\] you're given that \(a(t)=-32\), see if you can integrate above now..

OpenStudy (anonymous):

It is position. 500ft cliff above the ground=+500 Thrown upward with speed of 40 feet per second-it will eventually fall down to the ground which is negative=40t acceleration is negative because velocity is negative -32 feet per second squared t is seconds so second squared=-32t^2

OpenStudy (anonymous):

v(t) - 40 = (-32t^2)/2 ?

OpenStudy (anonymous):

Position s=0 means it is a at ground. You find t to know the time it took to reach the ground. You then used the velocity function by taking derivative of position. Plug that time it reached the ground to the velocity function.

ganeshie8 (ganeshie8):

nope, whats antiderivative of constant ?

OpenStudy (anonymous):

well you add a variable in this case t

OpenStudy (anonymous):

oh woops i see what i did wrong

OpenStudy (anonymous):

v(t) = -32t + 40: also @Shalante if we take your s(t) and we derive it until we get a(t) we get -64 and i think it should be -32

OpenStudy (anonymous):

no it is -64

OpenStudy (anonymous):

\[\int\limits v(t) = s(t) = -16t^2 + 40t + 500\] does this look right to you ?

OpenStudy (anonymous):

"se acceleration due to gravity as –32 feet per second squared"

OpenStudy (anonymous):

use *

OpenStudy (anonymous):

Actually, there lots of ways to do this. There is like 3

OpenStudy (anonymous):

yes that is right.

OpenStudy (anonymous):

i know but i have to show my work for this one so cant use physics equations sadly :(

OpenStudy (anonymous):

oh wait it is wrong. Supposed to be -32t^2-40t+500

OpenStudy (anonymous):

Make the position=0 by doing quadratic formula.

OpenStudy (anonymous):

use your calculator divide everything by -4 first

OpenStudy (empty):

Yeah as far as physics is concerned I think this would be pretty much the best way: \[K_i + V_i = K_f + V_f\] \[\frac{1}{2} m 40^2 + m 32*500 = \frac{1}{2} m v^2\] \[\sqrt{40^2 + 2*32*500} = v\]

ganeshie8 (ganeshie8):

careful, indefinite integral is unique only upto a constant

OpenStudy (anonymous):

interesting, i'm getting positive 40t because of the v(t) - v(0) part (when we added the v(0) to isolate the v(t))

ganeshie8 (ganeshie8):

\[s(t)-s(0)=\int\limits_0^t v(t) -16t^2 + 40t \] so you're right, you do get \(s(t) = -16t^2+40t+500\)

OpenStudy (anonymous):

okay so now just set s(t) = 0 to find when it touches the ground right?

OpenStudy (anonymous):

wait i thought s (t) = -16t^2 .... not -32t^2

ganeshie8 (ganeshie8):

physicists are overjoyed by the conservation of "anything", but i think math professors still want to see the integrals mess ;)

OpenStudy (anonymous):

Answer is negative because the ball's speed while hitting the ground is going downwards.

ganeshie8 (ganeshie8):

t=0 corresponds to s=500 right ?

ganeshie8 (ganeshie8):

recall that \(s(t)\) is the displacement from \(ground\) after \(t\) seconds

OpenStudy (anonymous):

\[-32t^2+40t+500=0 \] got \[t=3.376s,-4.626s\] Time should only be positive so use t=3.376s \[v(t)=-64t-40\] \[v(3.376)=-256.06 m/s\] are you sure this is right @ganeshie8

OpenStudy (anonymous):

shouldn't it be \[-16t^2 +40t + 500 = 0 \]

ganeshie8 (ganeshie8):

sry im not following the discussion between you and Jdosio... but you're not allowed to use direct formulas here right ?

OpenStudy (anonymous):

no only calculus

OpenStudy (anonymous):

okay my only positive root for v(t) is 6.978

OpenStudy (anonymous):

Ya use -16t^2

OpenStudy (anonymous):

so to find the velocity at that time we just plug that into the formula for v(t) right ?

OpenStudy (anonymous):

which gives me around -183.3030

OpenStudy (anonymous):

does that look about right to you guys?

OpenStudy (anonymous):

Nope it should be -40

OpenStudy (anonymous):

im using v(t) = -32t + 40 is that right?

OpenStudy (anonymous):

Answer is right, but the graph is wrong. Guess you do not need the graph.

OpenStudy (anonymous):

okay so then does everyone agree that the answer is -183.3030 !?

OpenStudy (anonymous):

Read it wrong at the beginning lol

OpenStudy (anonymous):

@Empty @ganeshie8 ?

OpenStudy (anonymous):

It is correct.

OpenStudy (anonymous):

okay ill stick around just incase anyone changes their mind, thanks for the help guys :D

OpenStudy (empty):

I got -183.3030 by doing it with the energy considerations alone, so good job! Same answer different route is a good indicator imo.

ganeshie8 (ganeshie8):

Yes the ball touches the ground after at `t=6.978` \(v(6.978) =40-32(6.978) \approx -183.303\) you want to avoid that negative sign though speed is \(183.303\) whatever unit

OpenStudy (empty):

Only reason why I think they'd want the negative sign there is because they gave -32 f/s^2 as the acceleration due to gravity, which imo is kind of awkward so I don't know what they want.

ganeshie8 (ganeshie8):

acceleration is a vector, so sign makes sense speed isn't

OpenStudy (anonymous):

ill just make it explicit that the velocity is downwards

ganeshie8 (ganeshie8):

they are not asking velocity

OpenStudy (empty):

Yeah you're right @ganeshie8 good point.

ganeshie8 (ganeshie8):

just persisting because, when you use the terms correctly, your professor knows that you understood the concepts.. that is for @Jdosio :)

OpenStudy (anonymous):

ill take note of that :)

OpenStudy (anonymous):

okay something weird has happened

OpenStudy (anonymous):

i found the speed of the ball at the time that the ball is at a height of 500 (the height of the cliff) and got -40 does that look oddly round to you guys ?

OpenStudy (anonymous):

@ganeshie8 @Empty @Shalante ?

OpenStudy (empty):

Well this is actually completely expected that the speed is the same for the same height, this is because it has no more potential or kinetic energy at this same height when you threw it. This is due to conservation of energy, which I realize is not exactly in the scope of what you're learning however it's clear at a glance to me because of this.

OpenStudy (anonymous):

okay so i should stick with my 183.3030 right?